OpenStudy (anonymous):

evaluate the integral from negative infinity to infinity of xe^(x^(2)) dx

OpenStudy (anonymous):

0 i believe

OpenStudy (anonymous):

yeah, its zero but i dont know how to evaluate it

OpenStudy (anonymous):

to get zero

OpenStudy (anonymous):

Let u=x^2 du/dx=2x du/2=x dx \[\int{xe^{x^2}}dx=1/2\int{e^u}du = \frac{e^u}{2}+C=\frac{e^{x^2}}{2}+C\] When x approaches infinity (or minus infinity, since it is squared), e^(x^2) approaches infinity as well. So the answer should be infinite. (Do you really mean e^(x^2) or is it maybe e^(-x^2)?)

OpenStudy (anonymous):

yeh it doesnt converge i remember learning about this forgot the reason though

OpenStudy (anonymous):

Intuitively, the function xe^(x^2) is odd, so the answer should be automatically zero.

OpenStudy (anonymous):

i has something to do with the limits... from -R to R it is 0 but something strange happens at infinity

OpenStudy (anonymous):

its: e^(-x^2)

OpenStudy (anonymous):

because e^(x^2) is obviously an even function, which is multiplied by the odd function x, and hence the function is odd. Odd functions always have that property.

OpenStudy (anonymous):

yes for finite values your right but not for infinite values

OpenStudy (anonymous):

thats a good way to look at it

OpenStudy (anonymous):

Oh! In that case, e^(-x^2) approaches zero for both positive infinity and negative infinity, so the answer is zero.

OpenStudy (anonymous):

oh ok :) thank u!

OpenStudy (anonymous):

no matter how good you get at maths it always comes down to signs to screw everything up